5. A company borrowed 560,000 by signing a 60-day, 5% note payable from its bank. Compute the total cash payment due on the note's maturity date.
Answer #1

5. Amount to be paid on maturity = Principal + Interest payment

Interest Payment = 560,000 x 5% x 60/365 = 4603

Cash payment at maturity = 560000 + 4603 = 564603